# COMPASS: Robust Feature Conformal Prediction for Medical Segmentation Metrics

### v1

@@ Acknowledgements
M.C. is supported by a fellowship from the Gulf Coast Consortia on the NLM Training Program in Biomedical Informatics and Data Science T15LM007093.
Code Availability
Code will be released at https://github.com/matthewyccheung/compass upon acceptance.
@@ Appendix A Proof of Theorem 1
Lemma 1 (Nestedness implies valid scores).
For any $x\in\mathcal{X}$, if $\{S_{\beta}(x)\}_{\beta\geq 0}$ is nested, then the nonconformity score
@@
By nestedness, the sets $S_{\beta}(x)$ expand monotonically in $\beta$. Thus, if $y$ is contained in some $S_{\beta^{\prime}}(x)$, then it is also contained in $S_{\beta}(x)$ for all $\beta\geq\beta^{\prime}$. Hence the infimum in the definition of $R(x,y)$ exists. Moreover, $y\in S_{\beta}(x)$ if and only if $\beta$ exceeds this infimum.
@@
Let $(X_{i},Y_{i})_{i=1}^{n+1}$ be exchangeable, and let $R_{i}:=R(X_{i},Y_{i})$. Then the rank of $R_{n+1}$ among $\{R_{1},\dots,R_{n},R_{n+1}\}$ is uniformly distributed on $\{1,\dots,n+1\}$, conditional on $D_{\mathrm{tr}}$.
Since the nonconformity score $R(\cdot,\cdot)$ is a measurable, deterministic function of $(X_{i},Y_{i})$ and $D_{\mathrm{tr}}$, the exchangeability of $(X_{i},Y_{i})$ implies exchangeability of $(R_{i})_{i=1}^{n+1}$. Therefore the rank of $R_{n+1}$ among all $n+1$ scores is uniform.
@@
By Lemma 1, $Y_{n+1}\in S_{\beta}(X_{n+1})$ if and only if $R_{n+1}\leq\beta$. Hence
@@
By Lemma 2, the rank of $R_{n+1}$ is uniform among $\{1,\dots,n+1\}$. Since $\hat{\beta}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th order statistic of $\{R_{1},\dots,R_{n}\}$, it follows that
$\mathbb{P}(R_{n+1}\leq\hat{\beta}\mid D_{\mathrm{tr}})\;\geq\;1-\alpha,$
and the claim follows.
Theorem 1 generalizes the standard split conformal method. In the classical setting, $S_{\beta}(x)=[-\beta,+\beta]$ corresponds to symmetric absolute-error balls around a predictor. Here, the sets $S_{\beta}(x)$ are generated by perturbations of latent features through $\Delta$ and evaluation under the downstream metric $h\circ g$. The proof requires only nestedness and exchangeability, so the same rank-uniformity argument ensures coverage. Thus, standard split conformal prediction is recovered as a special case of the present framework.
@@ Appendix B Asymmetric Perturbations
We extend the symmetric method to provide more flexible and efficient prediction intervals to the asymmetric case. This approach is particularly valuable when the relationship between latent perturbations and the downstream metric is non-linear or asymmetric. For instance, a positive perturbation of a certain magnitude might cause a large increase in volume, while a negative perturbation of the same magnitude results in only a small decrease. It also allows for independent control over the rate of under- and over-estimation errors, specified by $\alpha_{\mathrm{lo}}$ and $\alpha_{\mathrm{hi}}$. To account for this, we calculate two distinct non-conformity scores ($R_{lo}$$R_{hi}$) and two corresponding perturbation magnitudes ($\beta_{lo}$$\beta_{hi}$). By calibrating the upper and lower bounds independently, Asymmetric COMPASS constructs an interval that adapts to these skewed error distributions, often resulting in tighter bounds than its symmetric counterpart in such cases.
@@
Let the setup be the same as in the original theorem, and assume the family of sets $\{S_{\beta}(x)\}_{\beta\geq 0}$ is nested for every $x$ as per Definition 1.
@@
$R_{i,\mathrm{lo}}$ $:=\inf\{\beta\geq 0:m_{X_{i}}(-\beta)\leq Y_{i}\},$
$R_{i,\mathrm{hi}}$ $:=\inf\{\beta\geq 0:m_{X_{i}}(+\beta)\geq Y_{i}\}.$
@@
$S(X_{n+1}):=[m_{X_{n+1}}(-\hat{\beta}_{\mathrm{lo}}),\;m_{X_{n+1}}(+\hat{\beta}_{\mathrm{hi}})]$
@@ B.1 Proof of Theorem 3
Lemma 3 (Nestedness implies valid one-sided scores).
For any $x\in\mathcal{X}$, if $\{S_{\beta}(x)\}_{\beta\geq 0}$ is nested, then the nonconformity scores $R_{\mathrm{lo}}(x,y)$ and $R_{\mathrm{hi}}(x,y)$ are well-defined in $0,\infty]$ and satisfy
$m_{x}(-\beta)\leq y\quad$ $\Longleftrightarrow\quad R_{\mathrm{lo}}(x,y)\leq\beta,$
$m_{x}(+\beta)\geq y\quad$ $\Longleftrightarrow\quad R_{\mathrm{hi}}(x,y)\leq\beta.$
Nestedness implies that the lower bound $m_{x}(-\beta)$ is non-increasing in $\beta$ and the upper bound $m_{x}(+\beta)$ is non-decreasing in $\beta$. This one-sided monotonicity ensures that if a bound is satisfied for some $\beta^{\prime}$, it remains satisfied for all $\beta\geq\beta^{\prime}$. Hence, the infima in the definitions of the scores exist, and the equivalences hold.
@@
Let $(X_{i},Y_{i})_{i=1}^{n+1}$ be exchangeable. Let $R_{i,\mathrm{lo}}:=R_{\mathrm{lo}}(X_{i},Y_{i})$ and $R_{i,\mathrm{hi}}:=R_{\mathrm{hi}}(X_{i},Y_{i})$. Then the ranks of $R_{n+1,\mathrm{lo}}$ among $\{R_{1,\mathrm{lo}},\dots,R_{n+1,\mathrm{lo}}\}$ and $R_{n+1,\mathrm{hi}}$ among $\{R_{1,\mathrm{hi}},\dots,R_{n+1,\mathrm{hi}}\}$ are each uniformly distributed on $\{1,\dots,n+1\}$, conditional on $D_{\mathrm{tr}}$.
Since the nonconformity scores are measurable, deterministic functions of $(X_{i},Y_{i})$ and $D_{\mathrm{tr}}$, the exchangeability of the data pairs implies the exchangeability of each sequence of scores, $(R_{i,\mathrm{lo}})_{i=1}^{n+1}$ and $(R_{i,\mathrm{hi}})_{i=1}^{n+1}$. Therefore, the rank of the test score within each sequence is uniform.
@@
The total miscoverage event is the union of two one-sided miscoverage events. By the union bound,
$\mathbb{P}(Y_{n+1}\notin S(X_{n+1}))\leq\mathbb{P}(Y_{n+1}<m_{X_{n+1}}(-\hat{\beta}_{\mathrm{lo}}))+\mathbb{P}(Y_{n+1}>m_{X_{n+1}}(+\hat{\beta}_{\mathrm{hi}})).$
@@
This theorem provides a stronger guarantee than its symmetric counterpart. The proof relies on the same core principles of nestedness and exchangeability, but applies them to each bound independently before combining them with the union bound.
@@ Appendix C Algorithms
In the COMPASS framework, we decompose the model into an encoder $f$ that maps an input image $x$ to a representation $\hat{z}$, and a decoder $g$ that maps $\hat{z}$ to the final logits. A downstream metric function $h$ then maps the resulting segmentation to a scalar value. For COMPASS-L, the representation $\hat{z}$ is the final logit tensor itself; the encoder $f$ is the entire network and the decoder $g$ is the identity function. For COMPASS-J, the representation $\hat{z}$ is an internal feature map; the encoder $f$ is the part of the network up to that layer, and the decoder $g$ is the remainder of the network. For clarity in the algorithms, we define the perturbation and metric evaluation function $P(\hat{z},\Delta,\beta)$ as the interval $h(g(\hat{z}-\beta\Delta)),h(g(\hat{z}+\beta\Delta))]$.
C.1 Algorithm 1: COMPASS
With the COMPASS algorithm, we construct a valid prediction interval for a downstream metric. The interval is symmetric in the representation space, meaning we perturb the representation by the same magnitude, $\pm\hat{\beta}$, along a data-specific direction, $\Delta$. Following conventional CP procedures, we separate the algorithm into three stages: Training, Calibration, and Inference.
Algorithm 1 COMPASS
1:Training $D_{\mathrm{tr}}$, Calibration $D_{\mathrm{cal}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, Test $x_{n+1}$, $\alpha$, $L$.
2:Prediction interval $S_{\hat{\beta}}(x_{n+1})$.
@@ C.1 Algorithm 1: COMPASS
7:   $\mathcal{J}\leftarrow\{\sum_{\text{spatial}}\nabla_{\hat{z}_{j}}h(g(f(x_{j})))\}_{j\in D_{\mathrm{tr}}}$.
8:   $V_{L}\leftarrow\text{PCA}(\mathcal{J},L)$.

### v2

@@ Acknowledgements
MC would like to acknowledge support from a fellowship from the Gulf Coast Consortia on the NLM Training Program in Biomedical Informatics and Data Science T15LM007093.
Ethics Statement
We develop uncertainty quantification methods for medical image segmentation metrics using publicly available and anonymized datasets. The methods are intended as research tools to improve reliability and interpretability, but are not validated for clinical deployment. Risks include potential misinterpretation of uncertainty estimates. We therefore release code for research purposes only and encourage further evaluation with clinical experts before translation to practice.
Reproducibility Statement
We describe the algorithms in our paper extensively in Appendix C. We also describe the experimental details in Appendix E. Our code is available at https://github.com/matthewyccheung/compass. All packages used in the repository are publicly available. All datasets used are publicly accessible.
@@ Appendix A Proof of Theorem 1
Lemma 1 (Guaranteed Nestedness and Valid Scores).
For any $x\in\mathcal{X}$, let $m_{x}(b)$ be defined as in Theorem 1, and let
$S_{\beta}(x):=\left[\min_{b\in[-\beta,+\beta]}m_{x}(b),\max_{b\in[-\beta,+\beta]}m_{x}(b)\right].$
This family $\{S_{\beta}(x)\}_{\beta\geq 0}$ is nested. Furthermore, the nonconformity score
@@
Proof.
First, we prove nestedness. Let $0\leq\beta_{1}\leq\beta_{2}$. By definition, the perturbation interval $\beta_{1},+\beta_{1}]$ is a subset of $\beta_{2},+\beta_{2}]$. Let
$S_{\beta_{1}}(x)=\left[\min_{b\in[-\beta_{1},+\beta_{1}]}m_{x}(b),\max_{b\in[-\beta_{1},+\beta_{1}]}m_{x}(b)\right]$
and
$S_{\beta_{2}}(x)=\left[\min_{b\in[-\beta_{2},+\beta_{2}]}m_{x}(b),\max_{b\in[-\beta_{2},+\beta_{2}]}m_{x}(b)\right].$
The minimum of a function over a set is always greater than or equal to the minimum over a superset. Thus,
$\min_{b\in[-\beta_{1},+\beta_{1}]}m_{x}(b)\geq\min_{b\in[-\beta_{2},+\beta_{2}]}m_{x}(b).$
Conversely, the maximum over a set is less than or equal to the maximum over a superset. Thus,
$\max_{b\in[-\beta_{1},+\beta_{1}]}m_{x}(b)\leq\max_{b\in[-\beta_{2},+\beta_{2}]}m_{x}(b).$
Therefore, by definition of an interval, $S_{\beta_{1}}(x)\subseteq S_{\beta_{2}}(x)$, and the family is nested.
Second, we prove the equivalence $y\in S_{\beta}(x)\Longleftrightarrow R(x,y)\leq\beta$. Since we have just proven that the sets $S_{\beta}(x)$ are nested, they expand monotonically in $\beta$. If $y\in S_{\beta^{\prime}}(x)$ for some $\beta^{\prime}$, then $y\in S_{\beta}(x)$ for all $\beta\geq\beta^{\prime}$. The set $\{\beta\geq 0:y\in S_{\beta}(x)\}$ is therefore a closed interval of the form $R_{0},\infty)$ (or $0,\infty)$ if $y\in S_{0}(x)$). The infimum $R(x,y)$ is thus well-defined (as $R_{0}$), and $y\in S_{\beta}(x)$ if and only if $\beta\geq R(x,y)$. ∎
@@
Let $(X_{i},Y_{i})_{i=1}^{n+1}$ be exchangeable, and let $R_{i}:=R(X_{i},Y_{i})$ be derived from $D_{\mathrm{tr}}$. Then the rank of $R_{n+1}$ among $\{R_{1},\dots,R_{n},R_{n+1}\}$ is uniformly distributed on $\{1,\dots,n+1\}$, conditional on $D_{\mathrm{tr}}$.
Proof.
Since the nonconformity score $R(\cdot,\cdot)$ is a measurable, deterministic function of $(X_{i},Y_{i})$ and $D_{\mathrm{tr}}$, the exchangeability of the pairs $(X_{i},Y_{i})$ implies the exchangeability of the scalar scores $(R_{i})_{i=1}^{n+1}$. Therefore, by the standard properties of exchangeable random variables, the rank of $R_{n+1}$ among all $n+1$ scores is uniformly distributed on $\{1,\dots,n+1\}$. ∎
@@
Proof.
By Lemma 1, our construction of $S_{\beta}(x)$ guarantees nestedness and also guarantees that $Y_{n+1}\in S_{\beta}(X_{n+1})$ if and only if $R_{n+1}\leq\beta$. Applying this to our calibrated quantile $\hat{\beta}$, we have
@@
By Lemma 2, the rank of $R_{n+1}$ is uniform on $\{1,\dots,n+1\}$. By definition, $\hat{\beta}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value among the calibration scores $\{R_{1},\dots,R_{n}\}$. By the standard p-value argument for split conformal prediction, the probability that the test score $R_{n+1}$ is less than or equal to $\hat{\beta}$ is bounded.
$\mathbb{P}(R_{n+1}\leq\hat{\beta}\mid D_{\mathrm{tr}})\;\geq\;1-\alpha.$
Combining these, we have
$\mathbb{P}\!\left(Y_{n+1}\in S_{\hat{\beta}}(X_{n+1})\;\middle|\;D_{\mathrm{tr}}\right)=\mathbb{P}(R_{n+1}\leq\hat{\beta}\mid D_{\mathrm{tr}})\geq 1-\alpha,$
which proves the theorem. ∎
Connection to Standard CP.
Theorem 1 generalizes the standard split conformal method. In the classical setting for regression, $S_{\beta}(x)=[\hat{y}(x)-\beta,\hat{y}(x)+\beta]$ corresponds to symmetric absolute-error balls around a predictor $\hat{y}(x)$. This family is trivially nested. Here, the sets $S_{\beta}(x)$ are generated by perturbations of latent features through $\Delta$ and evaluation under the downstream metric $h\circ g$. By rigorously defining $S_{\beta}(x)$ as the $\min,\max]$ envelope, we guarantee the nestedness property required for the proof. The same rank-uniformity argument from exchangeability then ensures coverage. Thus, standard split conformal prediction is recovered as a special case of this framework.
@@ Appendix B Asymmetric Perturbations
We extend the symmetric method to provide more flexible and efficient prediction intervals to the asymmetric case. This approach is particularly valuable when the relationship between latent perturbations and the downstream metric is non-linear or asymmetric. For instance, a positive perturbation of a certain magnitude might cause a large increase in metric, while a negative perturbation of the same magnitude results in only a small decrease. It also allows for independent control over the rate of under- and over-estimation errors, specified by $\alpha_{\mathrm{lo}}$ and $\alpha_{\mathrm{hi}}$. To account for this, we calculate two distinct non-conformity scores ($R_{\mathrm{lo}}$, $R_{\mathrm{hi}}$) and two corresponding perturbation magnitudes ($\hat{\beta}_{\mathrm{lo}}$, $\hat{\beta}_{\mathrm{hi}}$). By calibrating the upper and lower bounds independently, Asymmetric COMPASS constructs an interval that adapts to these skewed error distributions, often resulting in tighter bounds than its symmetric counterpart in such cases.
@@
Let the setup be the same as in Theorem 1, with metric function $m_{x}(b):=(h\circ g)(\hat{z}(x)+b\Delta)$. We define the one-sided envelope functions:
$L_{x}(\beta)$ $:=\min_{b\in[-\beta,0]}m_{x}(b)$
$U_{x}(\beta)$ $:=\max_{b\in[0,+\beta]}m_{x}(b)$
These functions $L_{x}(\beta)$ and $U_{x}(\beta)$ are guaranteed to be non-increasing and non-decreasing in $\beta$, respectively.
@@
$R_{i,\mathrm{lo}}$ $:=\inf\{\beta\geq 0:L_{X_{i}}(\beta)\leq Y_{i}\},$
$R_{i,\mathrm{hi}}$ $:=\inf\{\beta\geq 0:U_{X_{i}}(\beta)\geq Y_{i}\}.$
@@
$S(X_{n+1}):=[L_{X_{n+1}}(\hat{\beta}_{\mathrm{lo}}),\;U_{X_{n+1}}(\hat{\beta}_{\mathrm{hi}})]$
@@ B.1 Proof of Theorem 3
Lemma 3 (Monotonic Bounds and Valid One-Sided Scores).
For any $x\in\mathcal{X}$, let $L_{x}(\beta)$ and $U_{x}(\beta)$ be defined as in Theorem 3. The function $L_{x}(\beta)$ is non-increasing in $\beta$, and the function $U_{x}(\beta)$ is non-decreasing in $\beta$. Furthermore, the nonconformity scores $R_{\mathrm{lo}}(x,y)$ and $R_{\mathrm{hi}}(x,y)$ are well-defined in $0,\infty]$ and satisfy
$L_{x}(\beta)\leq y\quad$ $\Longleftrightarrow\quad R_{\mathrm{lo}}(x,y)\leq\beta,$
$U_{x}(\beta)\geq y\quad$ $\Longleftrightarrow\quad R_{\mathrm{hi}}(x,y)\leq\beta.$
Proof.
Let $0\leq\beta_{1}\leq\beta_{2}$. By definition, the interval $\beta_{1},0]$ is a subset of $\beta_{2},0]$. The minimum of a function over a set is always greater than or equal to the minimum over a superset. Thus:
$L_{x}(\beta_{1})=\min_{b\in[-\beta_{1},0]}m_{x}(b)\geq\min_{b\in[-\beta_{2},0]}m_{x}(b)=L_{x}(\beta_{2}).$
This proves $L_{x}(\beta)$ is non-increasing. Similarly, $0,+\beta_{1}]$ is a subset of $0,+\beta_{2}]$. The maximum of a function over a set is always less than or equal to the maximum over a superset. Thus:
$U_{x}(\beta_{1})=\max_{b\in[0,+\beta_{1}]}m_{x}(b)\leq\max_{b\in[0,+\beta_{2}]}m_{x}(b)=U_{x}(\beta_{2}).$
This proves $U_{x}(\beta)$ is non-decreasing.
This one-sided monotonicity ensures that if a bound is satisfied for some $\beta^{\prime}$ (e.g., $L_{x}(\beta^{\prime})\leq y$), it remains satisfied for all $\beta\geq\beta^{\prime}$. Hence, the infima in the definitions of the scores exist, and the equivalences hold. ∎
@@
Let $(X_{i},Y_{i})_{i=1}^{n+1}$ be exchangeable. Let $R_{i,\mathrm{lo}}:=R_{\mathrm{lo}}(X_{i},Y_{i})$ and $R_{i,\mathrm{hi}}:=R_{\mathrm{hi}}(X_{i},Y_{i})$ be derived from $D_{\mathrm{tr}}$. Then the ranks of $R_{n+1,\mathrm{lo}}$ among $\{R_{1,\mathrm{lo}},\dots,R_{n+1,\mathrm{lo}}\}$ and $R_{n+1,\mathrm{hi}}$ among $\{R_{1,\mathrm{hi}},\dots,R_{n+1,\mathrm{hi}}\}$ are each uniformly distributed on $\{1,\dots,n+1\}$, conditional on $D_{\mathrm{tr}}$.
Proof.
Since the nonconformity scores are measurable, deterministic functions of $(X_{i},Y_{i})$ and $D_{\mathrm{tr}}$, the exchangeability of the data pairs implies the exchangeability of each sequence of scores, $(R_{i,\mathrm{lo}})_{i=1}^{n+1}$ and $(R_{i,\mathrm{hi}})_{i=1}^{n+1}$. Therefore, the rank of the test score within each sequence is uniform. ∎
@@
The total miscoverage event is the union of two one-sided miscoverage events:
$\mathbb{P}(Y_{n+1}\notin S(X_{n+1}))=\mathbb{P}(Y_{n+1}<L_{X_{n+1}}(\hat{\beta}_{\mathrm{lo}})\text{ or }Y_{n+1}>U_{X_{n+1}}(\hat{\beta}_{\mathrm{hi}})).$
By the union bound,
$\mathbb{P}(Y_{n+1}\notin S(X_{n+1}))\leq\mathbb{P}(Y_{n+1}<L_{X_{n+1}}(\hat{\beta}_{\mathrm{lo}}))+\mathbb{P}(Y_{n+1}>U_{X_{n+1}}(\hat{\beta}_{\mathrm{hi}})).$
@@
This theorem provides a rigorous guarantee for asymmetric intervals. The proof relies on the same core principles of exchangeability, but instead of assuming nestedness, we guarantee the required one-sided monotonicity by defining the bounds $L_{x}(\beta)$ and $U_{x}(\beta)$ as the $\min,\max]$ over the respective perturbation ranges.
@@ Appendix C Algorithms
In the COMPASS framework, we decompose the model into an encoder $f$ that maps an input image $x$ to a representation $\hat{z}$, and a decoder $g$ that maps $\hat{z}$ to the final segmentation. A downstream metric function $h$ then maps the segmentation to a scalar metric $\hat{y}$. For simplicity, we show the algorithm using the endpoint (and not the conservative envelope that requires a perturbation sweep during the calibration procedure) – i.e., we define the perturbation and metric evaluation function $P(\hat{z},\Delta,\beta)$ as the interval $h(g(\hat{z}-\beta\Delta)),h(g(\hat{z}+\beta\Delta))]$. This endpoint construction is computationally efficient and theoretically valid given the monotonic metric response observed in our experiments (Figures 2 and 12). The conservative envelope variant follows the same approach, but the symmetric binary search involves a perturbation sweep for each sample during the calibration step to find the minimum and maximum.
For COMPASS-L, the representation $\hat{z}$ is the logits; the encoder $f$ is the entire network, and the decoder $g$ is the identity function. The perturbation direction is simply $\Delta=\mathbf{1}$.
For COMPASS-J, the representation $\hat{z}$ is an internal feature map. To find the sensitive direction $\Delta$, we first compute the Jacobian of the metric with respect to the features, $)J=\nabla_{\hat{z}}h(g(\hat{z}$)). We then project this Jacobian onto the subspace spanned by the top $L$ principal components ($V_{L}$) of the training set Jacobians and reconstruct it. The final direction $\Delta$ is the normalized reconstruction: $\mathbf{d}=V_{L}V_{L}^{\top}J,\Delta=\frac{\mathbf{d}}{\|\mathbf{d}\|_{2}}$.
In general, the COMPASS framework follows several key steps:
  1. 1.
    Training: Assume an already trained base segmentation model, decomposed into an encoder $f$ and decoder $g$. If using COMPASS-J, compute the Jacobians of the metric with respect to the latent features $\hat{z}$ for the training set and perform PCA to identify the dominant sensitivity subspace $V_{L}$.
  2. 2.
    Monotonicity Verification: Perform a perturbation sweep on a subset of the data to verify that the metric response $m_{x}(b)$ is monotonic with respect to the perturbation magnitude along the computed direction $\Delta$. This empirical check justifies the use of the endpoints over the computationally expensive envelope construction.
  3. 3.
    Calibration: For each sample in the calibration set, compute the non-conformity score $R_{i}$ using a binary search. $R_{i}$ is the minimal perturbation magnitude $\beta$ such that the computed interval covers the ground truth. Finally, compute the calibrated quantile $\hat{\beta}$ from these scores using the standard finite-sample correction.
  4. 4.
[…]
C.1 Training
Algorithm 1 COMPASS Training
1:Training data $D_{\mathrm{tr}}$, $L$ (number of components for COMPASS-J).
2:Trained model $(f,g)$, and subspace $V_{L}$ (for COMPASS-J).
@@ C.1 Training
7:   Compute Jacobians: $\mathcal{J}\leftarrow\{\sum_{\text{spatial}}\nabla_{\hat{z}_{j}}h(g(f(x_{j})))\}_{j\in D_{\mathrm{tr}}}$.
8:   Compute PCA: $V_{L}\leftarrow\text{PCA}(\mathcal{J},L)$.
@@
C.2 Monotonicity Verification
Algorithm 2 Metric Monotonicity Verification
1:Model $(f,g,V_{L})$, Dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, $\beta_{\text{max}}$, $N_{\text{steps}}$.
2:Percentage of samples with monotonic metric response.
3:
4:procedure Verify-Monotonicity($D$, model, $h,\beta_{\text{max}},N_{\text{steps}}$)
5:  Initialize $count\leftarrow 0$.
6:  for $i=1,\dots,n$ do
7:   $\hat{z}_{i}\leftarrow f(x_{i})$
8:   if using COMPASS-J then
9:     $\mathbf{d}_{i}\leftarrow V_{L}V_{L}^{\top}\nabla_{\hat{z}}h(g(\hat{z}_{i}))$ $\triangleright$ Project Jacobian
10:     $\Delta_{i}\leftarrow\mathbf{d}_{i}/\|\mathbf{d}_{i}\|_{2}$ $\triangleright$ Normalize
11:   else
12:     $\Delta_{i}\leftarrow\mathbf{1}$
13:   end if
14:   $B\leftarrow\text{LinSpace}(-\beta_{\text{max}},\beta_{\text{max}},N_{\text{steps}})$. $\triangleright$ Discretized grid
15:   $M\leftarrow\emptyset$.
16:   for $b\in B$ do
17:     $M\leftarrow M\cup\{h(g(\hat{z}_{i}+b\Delta_{i}))\}$. $\triangleright$ Evaluate metric
18:   end for
19:   if Is-Sorted-Ascending($M$) then $\triangleright$ Check monotonicity
20:     $count\leftarrow count+1$.
21:   end if
22:  end for
23:  return $count/n$
24:end procedure
Algorithm 2 is a diagnostic step performed after training. It verifies that perturbing the features along the computed direction $\Delta$ results in a monotonic change in the metric (area). High monotonicity validates the use of the efficient endpoint-based interval construction in the subsequent calibration and inference steps.
C.3 Symmetric COMPASS
Algorithm 3 Symmetric COMPASS (Calibration & Inference)
1:Calibration $D_{\mathrm{cal}}$, Test $x_{n+1}$, $\alpha$, Model $(f,g,V_{L})$.
2:Prediction interval $S_{\hat{\beta}}(x_{n+1})$.
3:
4:procedure Calibration($D_{\mathrm{cal}}$, model, $h,\alpha$)
5:  Initialize scores $\mathcal{R}\leftarrow\emptyset$.
6:  for $i=1,\dots,n$ do
7:   Define $\hat{z}_{i}$ and compute $\Delta_{i}$ (as in Alg. 2).
8:   $R_{i}\leftarrow\textsc{Symmetric-Binary-Search}(\hat{z}_{i},\Delta_{i},y_{i})$.
9:   $\mathcal{R}\leftarrow\mathcal{R}\cup\{R_{i}\}$.
10:  end for
11:  $\hat{\beta}\leftarrow\text{Quantile}(\mathcal{R},\frac{\lceil(1-\alpha)(n+1)\rceil}{n})$.
12:  return $\hat{\beta}$
13:end procedure